# Agentic Program Verification

### v1

@@ Abstract.
In this work, we present a first LLM agent, AutoRocq, for conducting program verification. Unlike past works, which rely on extensive training of LLMs on proof examples, our agent learns on-the-fly and improves the proof via an iterative refinement loop. The iterative improvement of the proof is achieved by the proof agent communicating with the Rocq (formerly Coq) theorem prover to get additional context and feedback. The final result of the iteration is a proof derivation checked by the Rocq theorem prover. In this way, our proof construction involves autonomous collaboration between the proof agent and the theorem prover. This autonomy facilitates the search for proofs and decision-making in deciding on the structure of the proof tree.
Experimental evaluation on SV-COMP benchmarks and on Linux kernel modules shows promising efficacy in achieving automated program verification. As automation in code generation becomes more widespread, we posit that our proof agent can be potentially integrated with AI coding agents to achieve a generate and validate loop, thus moving closer to the vision of trusted automatic programming.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Under Review; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
The widespread adoption of AI-generated code has transformed the landscape of software development. Today, more than 25% of the new code at Google is generated by AI. Yet the increasing amounts of code often carry subtle semantic errors or vulnerabilities (pearce2025asleep) that may elude human review and testing, necessitating automated reasoning on program behaviors to engender trust. One promising approach to attain strong, machine-checkable evidence about program properties is formal verification, which has been pivotal in certifying high-assurance systems for decades, such as microkernels (DBLP:conf/sosp/KleinEHACDEEKNSTW09), compilers (leroy2016compcert), and databases (verified_db).
@@
Formal verification techniques commonly (1) reduce programs and specifications to logical formulas as proof obligations, and (2) discharge those obligations to automatic or interactive theorem provers (ITPs). Common ITPs, such as Isabelle/HOL and Rocq/Coq, automatically reduce the proof goals by applying user-supplied proof steps called tactics. Unfortunately, formal verification sees limited adoption in production systems, as it is an extremely time- and skill-intensive undertaking. Specifically, (1) capturing program behaviors formally requires significant efforts in manually annotating specifications and crafting loop invariants, and (2) generating a mathematically rigorous proof to validate verification conditions demands a high degree of expertise. For instance, the verification of the seL4 microkernel (DBLP:conf/sosp/KleinEHACDEEKNSTW09) required 22 person-years of effort, while the proofs for the CompCert compiler took 6 person-years and 100,000 lines of Rocq code—eight times longer than the implementation itself (leroy2016compcert). Most of the specifications and proofs are meticulously crafted by the developers. Recently, advances in Large Language Model (LLM) agents have been transformative across various disciplines. In particular, LLMs have demonstrated remarkable capabilities in both code understanding (ahmad2021unified) and general mathematical reasoning (alpha-proof)—two foundational capabilities required by program verification. In this work, we thus ask the following question: Can LLM agents be leveraged for automatic and end-to-end verification of real-world programs?
@@
State-of-the-art and what is needed. Since there have been approaches to discuss proof automation by leveraging LLMs (lu2025palm; rango-icse25) and other machine-learning-based techniques (yang2019coqgym; sanchez2024qedcartographer; sanchez2020generating)—let us try to understand the need for a new approach. First of all, it is reasonable to use machine learning models to learn and predict individual proof steps or tactics in a proof. This has been accomplished in several works, including Proverbot9001 (sanchez2020generating). It is also possible for LLMs to generate whole proofs, albeit potentially incorrect ones, and then have a repair step to correct them, if possible—an approach studied in PALM (lu2025palm). Rango (rango-icse25) goes one step further. Given a sequence of proof steps, it uses an ITP for yes/no validation of the proof steps, thereby allowing the trial of different tactics.
@@
Evaluation. We thoroughly evaluate our approach on existing mathematical lemmas (yang2019coqgym), as well as lemmas systematically extracted from SV-COMP programs (svcomp). The SV-COMP programs’ lemmas capture intricate code logic and properties, which are more representative of program verification. Evaluation shows that AutoRocq significantly outperforms state-of-the-art approaches. Specifically, AutoRocq is capable of proving 51.1% mathematical lemmas and 30.9% program lemmas, exceeding baseline approaches by 20.8% to 343.0% on mathematical lemmas, and by 42.4% to 204.6% on program lemmas. Among the successes, 142 lemmas (98 mathematical lemmas and 44 program lemmas) are uniquely proved by AutoRocq, due to its agentic design and access to proof contexts. We also conduct a case study to verify the code in Linux kernel modules (i.e., memory management utilities), where AutoRocq automatically verifies 12 lemmas for the function correctness property, compared to only 2–10 lemmas verified by baseline approaches.
@@
  • We build the first proof automation agent, AutoRocq, which is highly effective in automatic proof generation. It acts as an interpreter of proof representations and actively collaborates with the Rocq prover to construct proofs.
@@
  • We conduct an empirical study to demonstrate proof context- and structure-awareness can help LLM agents reason about software formally.
@@ 2.1. Formal Program Verification
Formal program verification uses mathematically rigorous methods to prove that a program satisfies its specifications. Unlike traditional testing, which can only expose errors for specific inputs or executions, formal verification provides logical guarantees that a program behaves correctly for all possible executions within a given semantic model. Within this landscape, a prominent paradigm is deductive verification. In this approach, the program code is annotated with formal specifications, such as preconditions, postconditions, and loop invariants. These annotations form formal contracts that describe the program’s intended behavior at the module and function boundaries. The annotated code is then compiled into a finite set of logical proof obligations under an explicit semantic model. The discharge of all proof obligations implies that the program satisfies its specifications.
@@
To facilitate deductive verification, a variety of tools have been developed, including Dafny (DBLP:conf/lpar/Leino10; DBLP:journals/pacmse/MisuLM024), Verus (DBLP:journals/corr/abs-2303-05491), Viper (DBLP:conf/vmcai/0001SS16), and Frama-C (DBLP:journals/fac/KirchnerKPSY15). These tools support annotating code, generating proof obligations, and attempting to discharge them. Automated solvers (de2008z3; barbosa2022cvc5) can be used to handle simple proof obligations (e.g., quantifier-free linear arithmetic (ge2009complete)). More complex proof obligations—such as those involving non-linear reasoning, quantifier handling, or sophisticated inductive proofs—require human guidance. In such cases, interactive theorem provers (ITPs) are used, where auxiliary lemmas are introduced to capture intermediate properties, helping to structure and simplify the proofs of these obligations.
@@
Rocq (formerly Coq) proof assistant (coq1996coq) is a widely-used interactive theorem prover for semi-automated validation of residual proof obligations. In Rocq, proofs are constructed in a top-down manner from a proof goal—typically the statement of a theorem or lemma. Users iteratively apply proof tactics, which are simple or parameterized commands that decompose the goal into zero or more subgoals and guide the construction of a proof term. This proof term is type-checked by Rocq’s trusted kernel. A proof succeeds when the subgoal list becomes empty, and the output proof term is the formal certificate of validity. This certificate can be reproduced by replaying the exact sequence of tactics, referred to as the proof script. Conceptually, successful proof induces a proof tree rooted at the original proof goal. Each node represents a subgoal produced during the derivation, and each edge corresponds to a tactic application. The interactive workflow of Rocq supports incremental, exploratory proof development with frequent inspection and backtracking.
2.2. Machine Learning for Proof Generation
@@
Proof generation aims to automatically synthesize proof scripts that discharge the given proof obligations, which closes the last mile of program verification. Fueled by recent advances in AI, a growing body of work tackles this task with machine-learning methods. Early works (yang2019coqgym; sanchez2020generating; first2020tactok; sanchez2023passport) often cast this problem as a sequence generation task. They design smart representations of syntactic constraints (yang2019coqgym), proof states (first2020tactok), and fine-grained proof contexts (sanchez2023passport) to enable neural networks to generate valid tactic sequences. However, applying a tactic successfully does not always advance the proof, as it may yield equivalent or even harder subgoals. This mismatch limits standalone prediction networks, which fixate on next-step tactic generation without a reliable global progress signal. To address this, an orthogonal line of work (blaauwbroek2020tactician; sanchez2024qedcartographer) augments these networks with search strategies to guide tactic generation. For instance, QEDCartographer (sanchez2024qedcartographer) extends Proverbot9001 (sanchez2020generating) with a tactic selection model, trained through reinforcement learning with a fine-grained reward function.
@@
On the other hand, Large Language Models (LLMs), with their remarkable capabilities in mathematical reasoning and high-level understanding, have become a promising alternative. A recent study (lu2025palm) shows that LLMs can grasp the high-level structure of proofs, though they often stumble on rudimentary errors such as invalid references. Building on this observation, PALM first prompts the LLM to generate a complete proof script, and then applies deterministic repair mechanisms supplemented by the external tool CoqHammer (coqhammer) to patch errors. More recently, Rango (rango-icse25) fine-tuned an LLM as a knowledge base for tactic generation. At every step of the proof, Rango retrieves the most relevant proofs and lemmas for the current proof state based on the predefined similarity function, and then prompts its knowledge base to generate the next tactic.
These existing works have made significant advances in automated proof generation. However, most approaches rely on models that are trained or fine-tuned on large corpora of existing proofs to guide tactic generation. A widely-used large-scale training dataset is CoqGym (yang2019coqgym), which contains 71K human-written proofs that cover mathematics, computer hardware, and programming language. Models trained or fine-tuned on this dataset are effective in automatically generating proof scripts for mathematical theorems and libraries. Unfortunately, their effectiveness is reduced when applied to proofs about computer programs, which are typically more complex in both structure and semantics. Moreover, existing works adopt deterministic mechanisms to patch proof errors, retrieve relevant contexts, and generate subsequent tactics, preventing them from making adaptive, demand-driven decisions. For example, context retrieval is usually based solely on similarity functions, which may not provide sufficient information for generating the next tactic. Such a non-agentic context search strategy further limits the ability of these approaches.
@@ 3. Motivating Example
Although many automatic proving approaches have been proposed, to our knowledge, none of them prove lemmas in an agentic fashion. To motivate our agentic approach, AutoRocq, we use a concrete example shown in Figure 1. Figure 1(a) shows the proof obligation (wp_goal, lines 1-7) extracted from verifying the function correctness property of a real-world program cggmp2005b in SV-COMP (svcomp), along with the complete proof (lines 9-23) generated by AutoRocq. The proof tree representation that guides AutoRocq’s proving process is visualized in Figure 1(b). This example proof goal requires proving that “i1 = 10%Z ” (line 7) holds for integer-typed (Z) variables i1 and i under a complex set of hypotheses involving modular arithmetic, inequalities, and type constraints (lines 2-6). The hypotheses capture program semantics including path conditions (e.g., “i ¡= 10%Z ”) and non-overflow constraints (“(-2147483648%Z)%Z ¡= x ”), all discharged by Frama-C during automated verification (see Section 5.1.2 for more details). This example represents a typical lemma derived from real-world verification tasks. It captures intricate program logic and thus requires sophisticated reasoning about nested quantifiers, integer arithmetic, and logical constraints to handle. The proof tree in Figure 1(b) illustrates how AutoRocq systematically decomposes the problem through strategic tactic application, including case analysis with destruct that creates multiple proof branches, and maintains rich contextual information throughout the proving process. Such complex lemmas pose significant challenges to existing approaches (e.g., (rango-icse25; lu2025palm; sanchez2024qedcartographer; sanchez2020generating)), and none of them can prove this lemma during our experiments. We highlight these challenges to motivate AutoRocq’s agentic design in the following.
@@
Figure 1. (a) Proof obligation wp_goal extracted from benchmark52_polynomial in SV-COMP (svcomp) and its proof generated by AutoRocq, with (b) the proof tree constructed during proving, where lemmas highlighted are retrieved from the global context autonomously by AutoRocq.
Verbose and crowded context
@@
The local context of wp_goal carries rich semantic information—six hypotheses are present in the statement simultaneously. These hypotheses include conjunctions, disjunctions, inequalities, and type constraints. Conventional approaches, which mostly employ naive or similarity-based retrieval, could easily get a surfeit or an insufficiency of contextual information. For example, before proof generation, PALM (lu2025palm) simply retrieves all relevant premises from the global environment. In this case, it would be overwhelmed by hundreds of existing premises on integer arithmetic, most of which are irrelevant in the context. Conversely, Rango (rango-icse25) retrieves lemmas based on lexical similarities, an imprecise metric that often fails to identify semantically relevant lemmas. This is why Rango fails to prove the example: it cannot pinpoint the specific contextual assumptions on integer bounds and arithmetic properties required for the proof.
@@
Understanding of proving progress
A complicated lemma such as wp_goal requires a delicate sequence of tactics to prove. When generating such sequences, however, existing LLM-based methods lack structured representations of the proof process. They typically rely on simple goal lists rather than structured proof trees. This textual and linear representation of the proof state makes them overly focus on predicting a single tactic without a holistic view of the proving progress. As a result, they often fail to adapt to the evolving state of the partially generated proof. AutoRocq explicitly maintains a well-structured proof tree representation, as shown in Figure 1(b). It captures the hierarchical structure and dependencies of the proof derivation. The proof tree is built gradually as the proving advances. For instance, AutoRocq carefully decomposes the original goal into multiple subgoals with the destruct tactic, resulting in the branching in Figure 1. This structure-aware representation enables strategic reasoning about proof progress, tactical decisions, and effective tracking of proof dependencies. As such, our proof agent is able to interpret the proof derivation from a higher level, thus achieving more effective proof generation.
Harnessing the feedback
Developing proofs, especially for intricate lemmas, is a trial-and-error process, and interactive theorem provers are designed to provide timely feedback to guide the proving process. Nonetheless, existing approaches hardly leverage this opportunity to refine their tactics. For example, PALM (lu2025palm) only performs deterministic repairs to failed tactics, whereas Rango (rango-icse25) takes only binary signals from the proof assistant and retries the prediction again. AutoRocq implements an effective feedback mechanism that incorporates the feedback from the prover to refine individual tactics. In addition, it also recognizes persistent errors from the history, and conducts autonomous context searches to collect additional context progressively. This process allows our agent to learn continuously from both failures and successes. In this example, despite 48 failed tactic applications (59 attempts in total) during proving, AutoRocq eventually recovers and synthesizes a complete proof in a short time (i.e., 156.7 seconds).

### v3

# Agentic Verification of Software Systems

@@ Abstract.
In this work, we present a first LLM agent, AutoRocq, for conducting automatic verification of software systems. Unlike past works, which rely on extensive training of LLMs on proof examples, our agent learns on-the-fly and improves the proof via an iterative refinement loop. The iterative improvement of the proof is achieved by the proof agent communicating with the Rocq (formerly Coq) theorem prover to get additional context and feedback. The final result of the iteration is a proof derivation checked by the Rocq theorem prover. In this way, our proof construction involves autonomous collaboration between the proof agent and the theorem prover. This autonomy facilitates the search for proofs and decision-making in deciding on the structure of the proof tree. Experimental evaluation on SV-COMP benchmarks and on Linux kernel modules shows promising efficacy in achieving automated program verification. As automation in code generation becomes more widespread, we posit that our proof agent can be potentially integrated with AI coding agents to achieve a generate and validate loop, thus moving closer to the vision of trusted automatic programming.
Formal Methods; Program Verification; Theorem-proving; LLM Agent
copyright: ccdoi: 10.1145/3808164journalyear: 2026journal: PACMSEjournalvolume: 3journalnumber: FSEarticle: FSE157publicationmonth: 7submissionid: fse26mainb-p985-pccs: Software and its engineering Formal software verificationccs: Security and privacy Logic and verification
@@ 1. Introduction
The widespread adoption of AI-generated code has transformed the landscape of software development. Today, more than 25% of the new code at Google is generated by AI (Griffin, 2024). Yet the increasing amounts of code often carry subtle semantic errors or vulnerabilities (Pearce et al., 2025) that may elude human review and testing, necessitating automated reasoning on program behaviors to engender trust. One promising approach to attain strong, machine-checkable evidence about program properties is formal verification, which has been pivotal in certifying high-assurance systems for decades, such as microkernels (Klein et al., 2009), compilers (Leroy et al., 2016), and databases (Malecha et al., 2010).
@@
Formal verification techniques commonly (1) reduce programs and specifications to logical formulas as proof obligations, and (2) discharge those obligations to automatic or interactive theorem provers (ITPs). Common ITPs, such as Isabelle/HOL and Rocq/Coq, automatically reduce the proof goals by applying user-supplied proof steps called tactics. Unfortunately, formal verification sees limited adoption in production systems, as it is an extremely time- and skill-intensive undertaking. Specifically, (1) capturing program behaviors formally requires significant efforts in manually annotating specifications and crafting loop invariants, and (2) generating a mathematically rigorous proof to validate verification conditions demands a high degree of expertise. For instance, the verification of the seL4 microkernel (Klein et al., 2009) required 22 person-years of effort, while the proofs for the CompCert compiler took 6 person-years and 100,000 lines of Rocq code—eight times longer than the implementation itself (Leroy et al., 2016). Most of the specifications and proofs are meticulously crafted by the developers. Recently, advances in Large Language Model (LLM) agents have been transformative across various disciplines. In particular, LLMs have demonstrated remarkable capabilities in both code understanding (Ahmad et al., 2021) and general mathematical reasoning (Google DeepMind, 2024)—two foundational capabilities required by program verification. In this work, we thus ask the following question: Can LLM agents be leveraged for automatic and end-to-end verification of real-world programs?
@@
State-of-the-art and what is needed. Since there have been approaches to proof automation that leverage LLMs (Lu et al., 2024; Thompson et al., 2025; Thakur et al., 2024) and other machine-learning-based techniques (Yang and Deng, 2019; Sanchez-Stern et al., 2025, 2020) — let us understand the need for a new approach. First of all, it is reasonable to use machine learning models to learn and predict individual proof steps or tactics in a proof. This has been accomplished in several works, including Proverbot9001 (Sanchez-Stern et al., 2020). It is also possible for LLMs to generate whole proofs, albeit potentially incorrect ones, and then have a repair step to correct them, if possible—an approach studied in PALM (Lu et al., 2024). Rango (Thompson et al., 2025) goes one step further. Given a sequence of proof steps, it uses an ITP for yes/no validation of the proof steps, thereby allowing the trial of different tactics.
@@
Evaluation. We thoroughly evaluate our approach on existing mathematical lemmas (Yang and Deng, 2019; Thompson et al., 2025), as well as verification conditions systematically extracted from SV-COMP programs (Beyer and Strejček, 2025). These program-derived lemmas capture intricate code logic and properties, and are more representative of program verification tasks. We compare AutoRocq against five baselines, namely Rango (Thompson et al., 2025), PALM (Lu et al., 2024), Copra (Thakur et al., 2024), QEDCartographer (Sanchez-Stern et al., 2025), and Proverbot9001 (Sanchez-Stern et al., 2020), and results show that AutoRocq significantly outperforms these state-of-the-art approaches. Specifically, AutoRocq is capable of proving 48.0% mathematical lemmas and 30.9% verification lemmas, exceeding baseline approaches by 15.2% to 172.8% on mathematical lemmas, and by 10.6% to 204.6% on verification lemmas. Among the successes, 168 lemmas (140 mathematical lemmas and 28 verification lemmas) are uniquely proved by AutoRocq, due to its agentic design and access to proof contexts. We also conduct a case study to verify the code in Linux kernel modules (i.e., memory management utilities), where AutoRocq automatically verifies 12 lemmas for the function correctness property, compared to only 2–10 lemmas verified by baseline approaches.
@@
  • We build and make available [^1] The tool is released at https://github.com/NUS-Program-Verification/AutoRocq. the first proof automation agent, AutoRocq, which is highly effective in automatic proof generation. It acts as an interpreter of proof representations and actively collaborates with the Rocq prover to construct proofs.
@@
  • We conduct an empirical study to demonstrate that context- and structure-awareness can help LLM agents reason about software formally.
@@ 2.1. Formal Program Verification
Formal program verification uses mathematically rigorous methods to prove that a program satisfies its specifications. Unlike traditional testing, which only checks specific executions, formal verification provides logical guarantees for all executions within a given semantic model. A prominent paradigm is deductive verification, where a program is annotated with formal specifications such as pre/postconditions and loop invariants. These annotations define formal contracts for the intended behavior of the program, and the annotated code is translated into logical proof obligations. If all proof obligations are discharged, the program is verified to satisfy its specifications.
@@
To facilitate deductive verification, a variety of tools have been developed, including Dafny (Leino, 2010; Misu et al., 2024), Verus (Lattuada et al., 2023), Viper (Müller et al., 2016), and Frama-C (Kirchner et al., 2015). These tools support annotating code, generating proof obligations, and attempting to discharge them. Automated solvers (De Moura and Bjørner, 2008; Barbosa et al., 2022) can be used to handle simple proof obligations (e.g., quantifier-free linear arithmetic (Ge and De Moura, 2009)). More complex proof obligations—such as those involving non-linear reasoning, quantifier handling, or sophisticated inductive proofs—require human guidance. In such cases, interactive theorem provers (ITPs) are used, where auxiliary lemmas are introduced to capture intermediate properties, helping to structure and simplify the proofs of these obligations.
@@
Rocq (formerly Coq) proof assistant (The Coq Development Team, 2023) is a widely used interactive theorem prover for semi-automated validation of proof obligations. In Rocq, proofs are constructed top-down from a proof goal, i.e., the statement of a theorem or lemma. Users iteratively apply proof tactics, which transform the current goal into zero or more subgoals and guide the construction of a proof term. This proof term is then type-checked by Rocq’s trusted kernel. A proof succeeds when no subgoals remain, and the resulting proof term serves as the formal certificate of validity. The exact sequence of tactics, called the proof script, can be replayed to reproduce the proof. Conceptually, a successful proof induces a proof tree rooted at the original goal, where nodes are subgoals and edges are tactic applications. This interactive workflow supports incremental and exploratory proof development with frequent inspection and backtracking.
2.2. Machine Learning for Proof Automation
@@
Proof automation aims to automatically synthesize proof scripts that discharge proof obligations, addressing the last mile of program verification. Recent advances in AI have led to a growing body of learning-based approaches. Early works (Yang and Deng, 2019; Sanchez-Stern et al., 2020; First et al., 2020; Sanchez-Stern et al., 2023) often formulate this task as sequence generation, using representations of syntactic constraints (Yang and Deng, 2019), proof states (First et al., 2020), and fine-grained proof contexts (Sanchez-Stern et al., 2023) to generate tactic sequences. However, successfully applying a tactic does not always mean the proof has progressed, since it may produce equivalent or even harder subgoals. This limits standalone prediction models that focus only on next-step tactic generation without a reliable notion of global progress. To address this, another line of work (Blaauwbroek et al., 2020; Sanchez-Stern et al., 2025) combines tactic prediction with search strategies. For example, QEDCartographer (Sanchez-Stern et al., 2025) extends Proverbot9001 (Sanchez-Stern et al., 2020) with a reinforcement-learning-based tactic selection model.
@@
On the other hand, Large Language Models (LLMs) have emerged as a promising alternative due to their strong mathematical reasoning and high-level understanding. A recent study (Lu et al., 2024) shows that while LLMs can capture the high-level structure of proofs, they often make basic mistakes such as invalid references. Building on this observation, PALM first prompts the LLM to generate a complete proof script, and then applies deterministic repair mechanisms together with CoqHammer (Czajka and Kaliszyk, 2018) to fix errors. Copra (Thakur et al., 2024) adopts an in-context learning approach for Lean/Coq, iteratively proposing tactics, executing them in the ITP, and using error feedback to construct complete proofs under a query budget. More recently, Rango (Thompson et al., 2025) fine-tunes an LLM as a knowledge base for tactic generation. At each proof step, it retrieves relevant proofs and lemmas for the current proof state and then prompts the model to generate the next tactic.
These works have significantly advanced automated proof generation. However, most rely on models trained or fine-tuned on large proof corpora to guide tactic generation. A widely used dataset is CoqGym (Yang and Deng, 2019), which contains 71K human-written proofs spanning domains such as mathematics, hardware, and programming languages. While models trained on this dataset are effective for mathematical theorems and libraries, they are less effective for proofs about computer programs, which are often more complex in both structure and semantics (see Section 5.1.4).
@@ 3. Motivating Example
Although many automatic theorem-proving approaches have been proposed, to our knowledge, none of them prove lemmas in an agentic fashion. To motivate our agentic approach, AutoRocq, we use a concrete example shown in Figure 1. Figure 1(a) shows the proof obligation (wp_goal, lines 1-7) extracted from verifying the function correctness property of a real-world program cggmp2005b in SV-COMP (Beyer and Strejček, 2025), along with the complete proof (lines 9-23) generated by AutoRocq. The proof tree representation that guides AutoRocq’s proving process is visualized in Figure 1(b). This example proof goal requires proving that “i1 = 10%Z ” (line 7) holds for integer-typed (Z) variables i1 and i under a complex set of hypotheses involving modular arithmetic, inequalities, and type constraints (lines 2-6). The hypotheses capture program semantics including path conditions (e.g., “i ¡= 10%Z ”) and non-overflow constraints (“(-2147483648%Z)%Z ¡= x ”), all discharged by Frama-C during automated verification (see Section 5.1.3 for more details). This example represents a typical lemma derived from real-world verification tasks. It captures intricate program logic and thus requires sophisticated reasoning about nested quantifiers, integer arithmetic, and logical constraints to handle. The proof tree in Figure 1(b) illustrates how AutoRocq systematically decomposes the problem through strategic tactic application, including case analysis with destruct that creates multiple proof branches, and maintains rich contextual information throughout the proving process. Such complex lemmas pose significant challenges to existing approaches (e.g., (Thompson et al., 2025; Lu et al., 2024; Sanchez-Stern et al., 2025, 2020)), and none of them can prove this lemma during our experiments. We highlight these challenges to motivate AutoRocq’s agentic design in the following.
@@
Figure 1. (a) Proof obligation wp_goal extracted from benchmark52_polynomial in SV-COMP (Beyer and Strejček, 2025) and its proof generated by AutoRocq, with (b) the proof tree constructed during proving, where lemmas highlighted are retrieved from the global context autonomously by AutoRocq.
@@
Verbose and Crowded Context. The local context of wp_goal carries rich semantic information—six hypotheses are present in the statement simultaneously. These hypotheses include conjunctions, disjunctions, inequalities, and type constraints. Conventional approaches, which mostly employ naive or similarity-based retrieval, could easily get a surfeit or an insufficiency of contextual information. For example, before proof generation, PALM (Lu et al., 2024) simply retrieves all relevant premises from the global environment. In this case, it would be overwhelmed by hundreds of existing premises on integer arithmetic, most of which are irrelevant in the context. Conversely, Rango (Thompson et al., 2025) retrieves lemmas based on lexical similarities, an imprecise metric that often fails to identify semantically relevant lemmas. This is why Rango fails to prove the example: it cannot pinpoint the specific contextual assumptions on integer bounds and arithmetic properties required for the proof.
@@
Understanding of Proving Progress. A complicated lemma such as wp_goal requires a delicate sequence of tactics to prove. When generating such sequences, however, existing LLM-based methods lack structured representations of the proof process: they rely on simple goal lists rather than structured proof trees. This textual and linear representation of the proof state makes them overly focus on predicting a single tactic without a holistic view of the proving progress. As a result, they often fail to adapt to the evolving state of the partially generated proof. AutoRocq explicitly maintains a well-structured proof tree representation, as shown in Figure 1(b). It captures the hierarchical structure and dependencies of the proof derivation. The proof tree is built gradually as the proving advances. For instance, AutoRocq carefully decomposes the original goal into multiple subgoals with the destruct tactic, resulting in the branching in Figure 1. This structure-aware representation enables strategic reasoning about proof progress, tactical decisions, and effective tracking of proof dependencies. As such, our proof agent is able to interpret the proof derivation from a higher level, thus achieving more effective proof generation.
Harnessing the Feedback. Developing proofs, especially for intricate lemmas, is a trial-and-error process (Shi et al., 2025), and interactive theorem provers are designed to provide timely feedback to guide the proving process. Nonetheless, existing approaches hardly leverage this opportunity to refine their tactics. For example, PALM (Lu et al., 2024) only performs deterministic repairs to failed tactics, whereas Rango (Thompson et al., 2025) takes only binary signals from the proof assistant and retries the prediction again. AutoRocq implements an effective feedback mechanism that incorporates the feedback from the prover to refine individual tactics. In addition, it also recognizes persistent errors from the history, and conducts autonomous context searches to collect additional context progressively. This process allows our agent to learn continuously from both failures and successes. In this example, despite 48 failed tactic applications (59 attempts in total) during proving, AutoRocq eventually recovers and synthesizes a complete proof in a short time (i.e., 156.7 seconds).